OpenStudy (anonymous):

y + 3 = 2x y + 4 = 3x If you use the comparison method to solve the given system, what is the resulting equation?

OpenStudy (anonymous):

y+3=2x =>y=2x-3 and, y+4=3x =>y=3x-4 so comparing y values from two equations, 2x-3=3x-4 =>4-3=3x-2x =>1=x =>x=1 Putting the value of x in first equation, y+3=2(1) =>y=2-3=-1 so the results are, x=1, y=-1